# BMIKE-53: Investigating Cross-Lingual Knowledge Editing with In-Context Learning

### v2

@@ Abstract
This paper introduces BMIKE-53, a comprehensive benchmark for cross-lingual in-context knowledge editing (IKE) across 53 languages, unifying three knowledge editing (KE) datasets: zsRE, CounterFact, and WikiFactDiff. Cross-lingual KE, which requires knowledge edited in one language to generalize across others while preserving unrelated knowledge, remains underexplored. To address this gap, we systematically evaluate IKE under zero-shot, one-shot, and few-shot setups, incorporating tailored metric-specific demonstrations. Our findings reveal that model scale and demonstration alignment critically govern cross-lingual IKE efficacy, with larger models and tailored demonstrations significantly improving performance. Linguistic properties, particularly script type, strongly influence performance variation across languages, with non-Latin languages underperforming due to issues like language confusion.
@@
Ercong Nie1, 2  Bo Shao3   Mingyang Wang1,2,5  Zifeng Ding1,2,4  Helmut Schmid1   Hinrich Schütze1, 2 1LMU Munich, Germany   2Munich Center for Machine Learning, Germany 3Technical University of Munich, Germany   4University of Oxford, UK 5Bosch Center for Artificial Intelligence, Germany nie@cis.lmu.de   bo.shao@tum.de
@@ 1 Introduction
This study addresses the critical gap in cross-lingual knowledge editing by proposing an comprehensive and multidimensional investigation of in-context knowledge editing (IKE) methods. We introduce BMIKE-53, a multilingual benchmark spanning 53 languages and integrating three representative KE datasets: zsRE, which evaluates regular fact modifications; CounterFact, which examines counterfactual knowledge updates; and WikiFactDiff (WFD), which assesses real-world, temporally dynamic knowledge updates. This benchmark is the most comprehensive multilingual KE resource to date, unifying diverse KE datasets into a consistent format and expanding them into multiple languages using LLM-assisted translation. The wide linguistic coverage allows us to systematically analyze cross-lingual differences and their underlying causes.
@@
In summary, our contributions are as follows: i) We introduce BMIKE-53, the most comprehensive multilingual KE benchmark, covering 53 languages and three diverse KE datasets, which serves as a foundation for evaluating cross-lingual KE methods. ii) We extensively evaluate gradient-free cross-lingual KE methods under various IKE setups, providing valuable insights into the effectiveness of in-context learning for cross-lingual knowledge editing. iii) We conduct a detailed analysis of factors influencing cross-lingual KE performance, uncovering the impact of linguistic properties, script types, and language confusion on cross-lingual knowledge transfer.
@@ 2 Related Work
Traditional KE methods are primarily gradient-based. They typically introduce additional trainable parameters, such asMEND (Mitchell et al., 2021) and SERAC (Mitchell et al., 2022)–or edit specific parameters of the original model, as inROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). However, these methods have high computational demands and are difficult to scale. Recently, gradient-free methods, such as IKE (Zheng et al., 2023), MeLLo (Zhong et al., 2023), and ICE (Cohen et al., 2024), have been studied in KE for LLMs. Given the multilingual in-context learning capabilities of English-centric LLMs (Lai et al., 2023; Nie et al., 2024; Zhang et al., 2024a), the potential for cross-lingual KE appears promising. Recent cross-lingual KE work largely employs gradient-based methods (Xu et al., 2023; Wang et al., 2023; Beniwal et al., 2024; Wei et al., 2024). A notable gradient-free work is ReMaKE (Wang et al., 2024b), a cross-lingual retrieval-augmented KE method. However, their method is specifically applied to a rather special KE scenario – batch edit. In this setting, multiple knowledge pieces, such as the entire knowledge base, are edited simultaneously (Appendix  C).
@@ 5.2 Dataset-Specific Performance Patterns
We observe substantial cross-dataset variance in editing efficacy, especially with loc and port queries, indicative of inherent dataset complexity differences. While WikiFactDiff (WFD) achieves comparable reliability (rel) and generality (gen) scores to zsRE and CounterFact, it shows the lowest port performance. This discrepancy could be attributed to the real-world nature of WFD, where all knowledge–both the original and updated facts–is temporally recent. Portability queries in WFD require the model to reason over a second-order knowledge chain, where the correct answer depends on understanding the relationship between the updated fact and its broader context, CounterFact, on the other hand, benefits the most from metric-specific demonstrations, particularly for locality and portability. The compositional nature of CounterFact’s fabricated facts allows the model to leverage targeted demonstrations effectively.
@@ 5.4 Impact of Demonstration Strategies
The comparison of zero-shot, one-shot, few-shot mixed, and few-shot metric-specific setups reveals the importance of demonstration quality and quantity in cross-lingual IKE. Figure  4 illustrates how the performance of Llama3.1-8B evolves across setups for each query type and dataset. In the zero-shot setup, models rely solely on their pre-trained capabilities, resulting in limited performance. While one-shot setups provide modest improvements by familiarizing the model with the task format, they fail to address the challenges of loc. A single, randomly selected demonstration often confuses the model, particularly when the demonstration type does not align with the target query type. For loc queries, when encountering rel or gen demonstration types, this misalignment can lead the model to incorrectly repeat the edited knowledge in the target language, further degrading performance. Adding more demonstrations in the few-shot mixed setup yields limited performance gains, particularly for loc queries. However, when demonstrations are tailored to the specific query type being tested, as in the 8-shot metric-specific setup, notable gains are observed This setup achieves the highest scores for loc and portability, as shown in Figure  4. These findings underscore the importance of demonstration quality and specificity in maximizing the benefits of in-context learning for cross-lingual knowledge editing.
@@ 6 Language Performance Variance in Cross-Lingual IKE
To analyze cross-lingual performance variation, we use a normalized metric that uses the exact match (EM) of English as a reference to highlight cross-lingual transfer differences. Specifically, we calculate the ratio of each target language’s EM performance to English. As visually evident in Table  3, some languages perform particularly poorly, while others achieve results closer to English, motivating further investigation into the factors influencing these differences.
@@ 7 Conclusion
This research was conducted in accordance with the ACM Code of Ethics. The datasets used in this paper are publicly available. We do not intend to share any Personally Identifiable Data with this paper. Our project may raise awareness of these low-resource languages in the computational linguistics community.
@@ A.1 Language Coverage
BMIKE-53 encompasses a total of 53 languages. A comprehensive list of these languages can be found in Table  8. Additionally, the table outlines the linguistic feature similarities between each target language and English.
@@ A.4 Translation Quality Control
We used back-translation techniques to assess the quality of the translations. Specifically, each translated sentence was back-translated into English, We calculated the BLEU score and semantic similarity between the original and the back-translated English text. BLEU Score measures the formal similarity between the original and back-translated sentences. Semantic Similarity evaluates the semantic alignment between the original and back-translated sentences using cosine similarity in a sentence embedding space. The results of the back-translation evaluation are shown in Table  5.
@@ Appendix B BMIKE-53 with Other Baseline Methods
The primary purpose of this paper is to provide a basic gradient-free method for cross-lingual KE. However, it is helpful for better understanding our gradient-free method to compare it with other existing baseline methods, including gradient-based methods. Therefore, we also experimented on BMIKE-53 with commonly used baseline methods, including fine-tuning (FT), LoRA (Hu et al., ), ROME (Meng et al., 2022), and KN (Dai et al., 2022) in selected languages. We use the EasyEdit framework (Wang et al., 2024a) to conduct the baseline experiments. Table  6 shows the baseline method results.
@@ Appendix C Further Discussion on Related Work
As discussed in §2, a few previous studies have explored gradient-free knowledge editing methods, notably ReMaKE (Wang et al., 2024b). Our work diverges from these existing approaches in several key aspects, which we will elaborate on below. Regarding task setup, ReMaKE employs a cross-lingual retrieval-augmented strategy tailored for batch edits, allowing simultaneous modifications of multiple knowledge pieces, such as an entire knowledge base. Conversely, our approach focuses on individual knowledge edits, which do not involve cross-lingual retrieval. his fundamental difference sets our approach apart and addresses a unique aspect of knowledge editing. In prompt engineering, when designing the demonstrations of ICL, ReMaKE uses translation pairs of source and target language facts to make up the cross-lingual demonstrations, from which the model cannot learn real cross-lingual knowledge editing competencies like portability and locality. In contrast, our MIKE method provides four different types of cross-lingual ICL demonstrations. From these demonstrations, the model can effectively learn real cross-lingual knowledge editing.
@@ Appendix D Experiment Implementation Details
Table  7 displays the experiment implementation details. We use the models from HuggingFace (https://huggingface.co/meta-llama/Llama-3.2-3B, https://huggingface.co/meta-llama/Llama-3.1-8B).
Parameter Value
Model meta-llama/Llama-3.1-8B,
meta-llama/Llama-3.2-3B
Max. length 4096
Num. of demonstration 8
Type of demonstration Reliability, Generality,
Locality, Portability
Proportion of demo. 1:3:2:2
GPU Type NVIDIA A100-SXM4-80GB
Number of GPU 4
Running hours 72
Table 7: Experimental Implementation Details.
Appendix E Full Results
We show the full experimental results for all three tasks in Table  9 (zsRE), Table  10 (CounterFact), and Table  11 (WFD), respectively.
@@ Appendix E Full Results
be Belarusian Slavic 78.64 85.83 70.42 16.80 99.35
bg Bulgarian Slavic 85.78 85.83 68.38 13.73 99.01
@@
cs Czech Slavic 73.99 85.83 66.51 13.73 99.71
@@
hr Croatian Slavic 83.18 85.83 69.67 12.71 99.50
@@
pl Polish Slavic 78.64 85.83 65.29 15.03 99.63
@@
ru Russian Slavic 81.18 85.83 64.76 16.80 95.81
sk Slovak Slavic 82.16 85.83 70.66 15.03 99.55
sl Slovenian Slavic 80.59 85.83 75.58 15.03 99.62
@@
sr Serbian Slavic 79.60 85.83 72.94 12.71 99.23
@@
uk Ukrainian Slavic 84.73 85.83 74.38 15.03 99.28

### v3

@@ Abstract
This paper introduces BMIKE-53, a comprehensive benchmark for cross-lingual in-context knowledge editing (IKE) across 53 languages, unifying three knowledge editing (KE) datasets: zsRE, CounterFact, and WikiFactDiff. Cross-lingual KE, which requires knowledge edited in one language to generalize across others while preserving unrelated knowledge, remains underexplored. To address this gap, we systematically evaluate IKE under zero-shot, one-shot, and few-shot setups, incorporating tailored metric-specific demonstrations. Our findings reveal that model scale and demonstration alignment critically govern cross-lingual IKE efficacy, with larger models and tailored demonstrations significantly improving performance. Linguistic properties, particularly script type, strongly influence performance variation across languages, with non-Latin languages underperforming due to issues like language confusion. Code and data are publicly available at: https://github.com/ercong21/MultiKnow/.
@@
Ercong Nie1,2  Bo Shao3,6   Mingyang Wang1,2,4  Zifeng Ding1,2,5  Helmut Schmid1   Hinrich Schütze1,2 1LMU Munich, Germany   2Munich Center for Machine Learning, Germany 3Technical University of Munich, Germany   4Bosch Center for Artificial Intelligence, Germany 5University of Oxford, UK   6CISPA Helmholtz Center for Information Security, Germany nie@cis.lmu.de   bo.shao@tum.de
@@ 1 Introduction
This study addresses the critical gap in cross-lingual knowledge editing by proposing a comprehensive and multidimensional investigation of in-context knowledge editing (IKE) methods. We introduce BMIKE-53, a multilingual benchmark spanning 53 languages and integrating three representative KE datasets: zsRE, which evaluates regular fact modifications; CounterFact, which examines counterfactual knowledge updates; and WikiFactDiff (WFD), which assesses real-world, temporally dynamic knowledge updates. This benchmark is the most comprehensive multilingual KE resource to date, unifying diverse KE datasets into a consistent format and expanding them into multiple languages using LLM-assisted translation. The wide linguistic coverage allows us to analyze cross-lingual differences and their underlying causes systematically.
@@
In summary, our contributions are as follows:
i) We introduce BMIKE-53, the most comprehensive multilingual KE benchmark, covering 53 languages and three diverse KE datasets, which serves as a foundation for evaluating cross-lingual KE methods.
ii) We extensively evaluate gradient-free cross-lingual KE methods under various IKE setups, providing valuable insights into the effectiveness of in-context learning for cross-lingual knowledge editing.
iii) We conduct a detailed analysis of factors influencing cross-lingual KE performance, uncovering the impact of linguistic properties, script types, and language confusion on cross-lingual knowledge transfer.
@@ 2 Related Work
Traditional KE methods are primarily gradient-based. They typically introduce additional trainable parameters, such as MEND (Mitchell et al., 2021) and SERAC (Mitchell et al., 2022)—or edit specific parameters of the original model, as in ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). However, these methods have high computational demands and are difficult to scale. Recently, gradient-free methods, such as IKE (Zheng et al., 2023), MeLLo (Zhong et al., 2023), and ICE (Cohen et al., 2024), have been studied in KE for LLMs. Given the multilingual in-context learning capabilities of English-centric LLMs (Lai et al., 2023; Nie et al., 2024; Zhang et al., 2024a), the potential for cross-lingual KE appears promising. Recent cross-lingual KE work largely employs gradient-based methods (Xu et al., 2023; Wang et al., 2023; Beniwal et al., 2024; Wei et al., 2024). A notable gradient-free work is ReMaKE (Wang et al., 2024b), a cross-lingual retrieval-augmented KE method. However, their method is specifically applied to a rather special KE scenario—batch edit. In this setting, multiple knowledge pieces, such as the entire knowledge base, are edited simultaneously (Appendix  C).
@@ 5.2 Dataset-Specific Performance Patterns
We observe substantial cross-dataset variance in editing efficacy, especially with loc and port queries, indicative of inherent dataset complexity differences. While WikiFactDiff (WFD) achieves comparable reliability (rel) and generality (gen) scores to zsRE and CounterFact, it shows the lowest port performance. This discrepancy could be attributed to the real-world nature of WFD, where all knowledge–both the original and updated facts–is temporally recent. Portability queries in WFD require the model to reason over a second-order knowledge chain, where the correct answer depends on understanding the relationship between the updated fact and its broader context. CounterFact, on the other hand, benefits the most from metric-specific demonstrations, particularly for locality and portability. The compositional nature of CounterFact’s fabricated facts allows the model to leverage targeted demonstrations effectively.
@@ 5.4 Impact of Demonstration Strategies
The comparison of zero-shot, one-shot, few-shot mixed, and few-shot metric-specific setups reveals the importance of demonstration quality and quantity in cross-lingual IKE. Figure  4 illustrates how the performance of Llama3.1-8B evolves across setups for each query type and dataset. In the zero-shot setup, models rely solely on their pre-trained capabilities, resulting in limited performance. While one-shot setups provide modest improvements by familiarizing the model with the task format, they fail to address the challenges of loc. A single, randomly selected demonstration often confuses the model, particularly when the demonstration type does not align with the target query type. For loc queries, when encountering rel or gen demonstration types, this misalignment can lead the model to incorrectly repeat the edited knowledge in the target language, further degrading performance. Adding more demonstrations in the few-shot mixed setup yields limited performance gains, particularly for loc queries. However, when demonstrations are tailored to the specific query type being tested, as in the 8-shot metric-specific setup, notable gains are observed. This setup achieves the highest scores for loc and portability, as shown in Figure  4. These findings underscore the importance of demonstration quality and specificity in maximizing the benefits of in-context learning for cross-lingual knowledge editing.
@@ 6 Language Performance Variance in Cross-Lingual IKE
To analyze cross-lingual performance variation, we use a normalized metric that uses the exact match (EM) of English as a reference to highlight cross-lingual transfer differences. Specifically, we calculate the ratio of each target language’s EM performance to English. As is visually evident in Table  3, some languages perform particularly poorly, while others achieve results closer to English, motivating further investigation into the factors influencing these differences.
@@ 7 Conclusion
This research was conducted in accordance with the ACM Code of Ethics. The datasets used in this paper are publicly available. We do not intend to share any Personally Identifiable Data with this paper. Our project may raise awareness of these low-resource languages in the computational linguistics community. Artificial intelligence (AI) tools were utilized exclusively for language refinement during the preparation of this manuscript. The use of AI tools is disclosed here in accordance with current ethical guidelines and publishing policies.
Acknowledgments
We want to thank the anonymous reviewers for their valuable feedback and constructive suggestions, which have helped to improve the quality of this paper. This work was supported by the German Research Foundation (DFG) under grant SCHU 2246/14-1, Munich Center for Machine Learning (MCML), and China Scholarship Council (CSC).
@@ A.1 Language Coverage
BMIKE-53 encompasses a total of 53 languages. A comprehensive list of these languages can be found in Table  9. Additionally, the table outlines the linguistic feature similarities between each target language and English.
@@ A.4 Translation Quality Control
We used back-translation techniques to assess the quality of the translations. Specifically, each translated sentence was back-translated into English. We calculated the BLEU score and semantic similarity between the original and the back-translated English text. BLEU Score measures the formal similarity between the original and back-translated sentences. Semantic Similarity evaluates the semantic alignment between the original and back-translated sentences using cosine similarity in a sentence embedding space. The results of the back-translation evaluation are shown in Table  5.
@@ Appendix B BMIKE-53 with Other Baseline Methods
The primary purpose of this paper is to provide a basic gradient-free method for cross-lingual KE. However, it is helpful for a better understanding of our gradient-free method to compare it with other existing baseline methods, including gradient-based methods. Therefore, we also experimented on BMIKE-53 with commonly used baseline methods, including fine-tuning (FT), LoRA (Hu et al., ), ROME (Meng et al., 2022), and KN (Dai et al., 2022) in selected languages. We use the EasyEdit framework (Wang et al., 2024a) to conduct the baseline experiments. Table  7 shows the baseline method results.
Parameter Value
Model meta-llama/Llama-3.1-8B,
meta-llama/Llama-3.2-3B
Max. length 4096
Num. of demonstration 8
Type of demonstration Reliability, Generality,
Locality, Portability
Proportion of demo. 1:3:2:2
GPU Type NVIDIA A100-SXM4-80GB
Number of GPU 4
Running hours 72
Table 6: Experimental Implementation Details.
@@ Appendix C Further Discussion on Related Work
As discussed in §2, a few previous studies have explored gradient-free knowledge editing methods, notably ReMaKE (Wang et al., 2024b). Our work diverges from these existing approaches in several key aspects, which we will elaborate on below. Regarding task setup, ReMaKE employs a cross-lingual retrieval-augmented strategy tailored for batch edits, allowing simultaneous modifications of multiple knowledge pieces, such as an entire knowledge base. Conversely, our approach focuses on individual knowledge edits, which do not involve cross-lingual retrieval. This fundamental difference sets our approach apart and addresses a unique aspect of knowledge editing. In prompt engineering, when designing the demonstrations of ICL, ReMaKE uses translation pairs of source and target language facts to make up the cross-lingual demonstrations, from which the model cannot learn real cross-lingual knowledge editing competencies like portability and locality. In contrast, our MIKE method provides four different types of cross-lingual ICL demonstrations. From these demonstrations, the model can effectively learn real cross-lingual knowledge editing.
@@ Appendix D Experiment Implementation Details
Table  6 displays the experiment implementation details. We use the models from HuggingFace (https://huggingface.co/meta-llama/Llama-3.2-3B, https://huggingface.co/meta-llama/Llama-3.1-8B).
Appendix E Additional Experiments with Different Models
To investigate the generalization of BMIKE-53, we also benchmark several more multilingual and English-centric LLMs, including Bloomz-7b1  [^2] https://huggingface.co/bigscience/bloomz-7b1, Ministral-8B  [^3] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410, and Qwen2.5-7B  [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct. We use a subset of 10 languages in BMIKE-53, and report the average results across the following 10 languages: de, es, fa, fr, hu, ja, ko, nl, ru, tr, vi, zh. Table  8 shows the results of the additional experiments. For better comparison, we also report the average results across the 10 languages for Llama3.1-8B and Llama3.2-3B.
zsre CounterFact WikiFactDiff
rel gen loc port rel gen loc port rel gen loc port
Bloomz-7b1 zero-shot 27.13 26.53 1.85 4.40 24.78 23.97 5.53 6.62 28.68 28.42 3.00 1.56
one-shot 28.71 27.94 1.06 3.57 26.21 25.40 4.84 5.59 24.54 24.28 2.10 1.10
8-shot (mix) 31.26 30.86 2.42 5.87 31.86 31.33 8.29 6.97 26.87 26.78 4.13 2.48
8-shot (metric) 33.83 33.43 2.35 6.20 34.85 34.43 11.94 8.93 32.50 32.45 4.62 3.46
Ministral-7b zero-shot 50.63 49.30 8.62 12.16 41.65 39.97 16.24 13.89 51.48 48.97 5.71 3.06
one-shot 73.21 72.76 12.74 21.40 77.69 76.27 14.68 26.22 53.54 52.71 4.17 2.86
8-shot (mix) 73.49 73.06 14.97 28.62 68.55 66.54 20.77 30.67 54.71 53.80 5.84 7.31
8-shot (metric) 73.48 73.44 15.38 31.77 75.61 74.12 42.31 43.88 59.58 58.89 9.31 11.78
Qwen2.5-7b zero-shot 41.76 41.32 2.92 8.82 40.23 38.12 3.98 11.28 43.08 39.07 3.36 2.62
one-shot 60.27 59.15 1.57 11.51 65.28 62.06 8.54 16.35 54.89 50.67 3.76 2.44
8-shot (mix) 63.36 63.05 3.51 16.51 64.24 62.05 15.39 21.34 53.48 51.11 4.69 3.63
8-shot (metric) 63.71 63.03 9.73 20.01 66.30 64.24 35.24 29.98 58.52 57.58 7.65 5.83
Table 8: Additional experimental results with more various models. Average results across 10 languages (de, es, fa, fr, hu, ja, ko, nl, ru, tr, vi, zh).
Appendix F Full Results
We show the full experimental results for all three tasks in Table  10 (zsRE), Table  11 (CounterFact), and Table  12 (WFD), respectively.
@@ Appendix F Full Results
be Belarusian Indo-European (Slavic) 78.64 85.83 70.42 16.80 99.35
bg Bulgarian Indo-European (Slavic) 85.78 85.83 68.38 13.73 99.01
@@
cs Czech Indo-European (Slavic) 73.99 85.83 66.51 13.73 99.71
@@
hr Croatian Indo-European (Slavic) 83.18 85.83 69.67 12.71 99.50
@@
pl Polish Indo-European (Slavic) 78.64 85.83 65.29 15.03 99.63
@@
ru Russian Indo-European (Slavic) 81.18 85.83 64.76 16.80 95.81
sk Slovak Indo-European (Slavic) 82.16 85.83 70.66 15.03 99.55
sl Slovenian Indo-European (Slavic) 80.59 85.83 75.58 15.03 99.62
@@
sr Serbian Indo-European (Slavic) 79.60 85.83 72.94 12.71 99.23
@@
uk Ukrainian Indo-European (Slavic) 84.73 85.83 74.38 15.03 99.28